Suppose Bob owns 2,000 shares of Company X and 10,000 shares of Company Y. The total value of Bob's holdings of these two companies is $372,000.
Suppose Frank owns 8,000 shares of Company X and 6,000 shares of Company Y. The total value of Franks holdings of these two companies is $400,000.
a. Write equations for Bob and Frank's holdings. Use the variables X and Y to represent the values of shares of Company X and Company Y.
b. Solve for the value of a share of Company X and Company Y. Show your work so you can get partial credit even if your final answer is wrong.

Answer by jrfrunner(365)

a. Write equations for Bob and Frank's holdings. Use the variables X and Y to represent the values of shares of Company X and Company Y.
-
Bob: 2000*x+10000*y=372000
Frank: 8000*x+6000*y=400000
===
b. Solve for the value of a share of Company X and Company Y.
1) simplify the equations for Bob and Frank's holdings
2000*x+10000*y=372000 divide both sides by 2000
x+5y=186
8000*x+6000*y=400000 divide both sides by 2000
4x+3y=200
--
your new equations are
x+5y=186
4x+3y=200
--
multiply first equation by -4
-4x-20y=-744
4x+3y=200
add both equations
--
-17y=-544
y=32
substitute this into x+5y=186
x=186-5(32)=26
answer x=26, y=32
--
validate
subtitute these value into original equations
Bob: 2000*x+10000*y=372000---> 2000*26+10000*32 does this = 372000??
Frank: 8000*x+6000*y=400000----> 8000*26+6000*32 does this = 400000